Question 1 Two charges are set on the x-axis 15.6 cm away from each other. The charges are 16.4 nC and-6.90 nC. Calculate the electric field at the point on the x-axis between the charges where the electric potential due to these two charges is zerdo 4.13x104 N/C -6.40x 104 N/C 878 N/C ● 7.23x103 N/C
Answer #1

Let x is the point from 16.4nC charge where total potential will be zero due to these charges.

Electrical Potential = kq/r

Electrical Potential due to charge 16.4nC = 9×10^9 × 16.4×10^-9/x = 147.6/x

Electrical potential due to charge -6.9nC = -6.9×9/(15.6-x) = -62.1/(15.6-x)

Total electrical potential = 147.6/x - 62.1/(15.6-x) = 0

37.07 - 2.3768x = x

x = 10.977 cm

Electrical field due to these charges at that point = 147.6/(10.977×10^-2)^2 + 62.1/(4.622×10^-2)^2

Electrical field = 4.13×10^4 N/C